OpenStudy (anonymous):

Solve the equation 4x^2-45x+225=15x+7. Is there a formula im supposed to use? How do i start

OpenStudy (anonymous):

subtract 15x and 7 from both sides so you have it set = 0

OpenStudy (anonymous):

\[4x^2-45x+225=15x+7\] \[4x^2-60x+218=0\]

OpenStudy (anonymous):

and do the quadratic formula after?

OpenStudy (anonymous):

now maybe divide by 2 and get \[2x^2-30x+119=0\] will make solving easier. yes, then the quadratic formula